# On the accuracy of posterior recovery with neural network emulators

### v1

@@ Appendix B Emulator Bias
In fig. 7 we show the ‘emulator bias’ as defined by eq. 37 for each parameter and each noise level. The average bias increases with decreasing noise level, as we would expect, from 0.04 for 250 mK to 0.33 at 5 mK. The maximum bias at $\sigma=5$ mK is 0.69 which is less than the threshold for accurate posterior recovery defined in DJ23 of 1.
Figure 7: The plot shows the Emulator Bias, as defined in eq. 37, for each parameter at each noise level. DJ23 suggests that an emulator bias $>1$ indicates a poor recovery of the posterior. DJ23 finds that for $\sigma\leq 25$ mK the emulator bias begins to exceed this limit for $\log T_{\mathrm{min}}$ and $\gamma_{\mathrm{lo}}$ in particular. While we find that the emulator bias increases with decreasing noise, on average, we find that it never exceeds a value of 1 for any of the parameters.
@@ Appendix C Reproducibility of results
In our analysis, we use the default number of live points recommended in the polychord implementation of nested sampling, corresponding to $n_{\mathrm{live}}=200$ or 25 per dimension. We rerun our analysis on the fiducial mock data with 25 mK to check that $n_{\mathrm{live}}$ is high enough when using both the emulator and ARES. The results are shown in fig. 8. We find that 200 live points is enough to recover consistent chains on repeated runs and report the corresponding Bayesian evidences in table 4.
@@
Run 1 $-2297.18\pm 0.25$ $-2297.01\pm 0.25$
Run 2 $-2297.26\pm 0.25$ $-2296.94\pm 0.25$
Table 4: The Bayesian evidence for the four fits shown in fig. 8. We repeat the inference on the fiducial ARES signal with 25 mK noise with both globalemu and ARES to check that the sampler, polychord, is set up appropriately such that we recover consistent chains. fig. 8 shows that the posteriors are visually the same and the table here shows that the evidences are consistent.
Figure 8: Left Panel: We show two sets of chains from running polychord with $n_{\mathrm{live}}=200$ on the globalemu fit to the fiducial ARES signal with $\sigma=25$ mK. The chains are visually similar and the Bayesian evidences for both fits are consistent within the sampling error as shown in table 4. Right Panel: The same as the left panel, but fitting the fiducial ARES signal plus noise directly with ARES. Again the posteriors appear consistent, and the evidences are consistent with error.

### v2

@@ Appendix B Emulator Bias
In fig.˜7 we show the ‘emulator bias’ as defined by eq.˜37 for each parameter and each noise level. The average bias increases with decreasing noise level, as we would expect, from 0.04 for 250 mK to 0.33 at 5 mK. The maximum bias at $\sigma=5$ mK is 0.69 which is less than the threshold for accurate posterior recovery defined in DJ23 of 1.
Figure 7: The plot shows the Emulator Bias, as defined in eq.˜37, for each parameter at each noise level. DJ23 suggests that an emulator bias >1>1 indicates a poor recovery of the posterior. DJ23 finds that for $\sigma\leq 25$ mK the emulator bias begins to exceed this limit for $\log T_{\mathrm{min}}$ and $\gamma_{\mathrm{lo}}$ in particular. While we find that the emulator bias increases with decreasing noise, on average, we find that it never exceeds a value of 1 for any of the parameters.
@@ Appendix C Reproducibility of results
In our analysis, we use the default number of live points recommended in the polychord implementation of nested sampling, corresponding to $n_{\mathrm{live}}=200$ or 25 per dimension. We rerun our analysis on the fiducial mock data with 25 mK to check that $n_{\mathrm{live}}$ is high enough when using both the emulator and ARES. The results are shown in fig.˜8. We find that 200 live points is enough to recover consistent chains on repeated runs and report the corresponding Bayesian evidences in table˜4.
@@
Run 1 $297.18\pm 0.25 22$$297.01\pm 0.25$
Run 2 $297.26\pm 0.25 22$$296.94\pm 0.25$
Table 4: The Bayesian evidence for the four fits shown in fig.˜8. We repeat the inference on the fiducial ARES signal with 25 mK noise with both globalemu and ARES to check that the sampler, polychord, is set up appropriately such that we recover consistent chains. fig.˜8 shows that the posteriors are visually the same and the table here shows that the evidences are consistent.
Figure 8: Left Panel: We show two sets of chains from running polychord with $n_{\mathrm{live}}=200$ on the globalemu fit to the fiducial ARES signal with $\sigma=25$ mK. The chains are visually similar and the Bayesian evidences for both fits are consistent within the sampling error as shown in table˜4. Right Panel: The same as the left panel, but fitting the fiducial ARES signal plus noise directly with ARES. Again the posteriors appear consistent, and the evidences are consistent with error.